# Modelling Global Trade with Optimal Transport

### v2

@@ Abstract
Global trade is shaped by a complex mix of factors beyond supply and demand, including tangible variables like transport costs and tariffs, as well as less quantifiable influences such as political and economic relations. Traditionally, economists model trade using gravity models, which rely on explicit covariates but often struggle to capture these subtler drivers of trade. In this work, we employ optimal transport and a deep neural network to learn a time-dependent cost function from data, without imposing a specific functional form. This approach consistently outperforms traditional gravity models in accuracy while providing natural uncertainty quantification. Applying our framework to global food and agricultural trade, we show that the global South suffered disproportionately from the war in Ukraine’s impact on wheat markets. We also analyze the effects of free-trade agreements and trade disputes with China, as well as Brexit’s impact on British trade with Europe, uncovering hidden patterns that trade volumes alone cannot reveal.
@@
1 Department of Applied Mathematics and Theoretical Physics, University of Cambridge, Cambridge CB3 0WA, United Kingdom; 2 Department of Mathematics, Imperial College London, London SW7 2AZ, United Kingdom; 3 Mathematics Institute, University of Warwick, Coventry CV4 7AL, UK; 4 The Alan Turing Institute, London NW1 2DB, United Kingdom; 5 School of Business and Management, Queen Mary University of London, London E1 4NS, United Kingdom
@@ Introduction
[…]
[lines=2,lraise=0.05,findent=0.1em, nindent=0em]International trade serves as the backbone of the world economy, distributing goods and connecting markets through global logistics networks. Its dynamics are driven by numerous factors beyond mere supply and demand, such as tariffs, non-tariff policy barriers, political and economic tensions, and disruptions caused by accidents, conflicts, and civil wars. Among all traded commodities, agricultural and food products hold particular interest for policymakers and the general public due to their significant volume, high trade value, and critical role in food security and resilience [1, 2]. Consumer food prices are a product of all the complexly interwoven factors governing trade. However, they do not always directly reflect the ease of doing business between any two countries. For instance, in May 2020, China imposed an 80% tariff on Australian barley, leading to a major restructuring of global supply chains (see fig. LABEL:fig:Barley_trade): Chinese demand was suddenly met from France, Canada, and Argentina, while Australia started exporting surplus barley e.g. to Saudi Arabia. Despite these shifts, for the next five months the global barley price barely budged [3, 4].
Modelling global trade has garnered significant attention in the economic literature, with gravity models being the most widely used approach [5, 6, 7, 8, 9]. These models, named for their direct analogy to Newton’s law of gravity, assume that the total trade $T_{ij}$ of a given commodity between two countries $i$ and $j$ is proportional to the total output $O_{i}$ of the source country and the total expenditure $E_{j}$ of the destination country, as well as inversely related to a ‘distance’ between them:
@@
where $\pi_{i}$ and $\chi_{j}$ are exogenous exporter and importer-side regressors [10], $\rho_{ij}$ are bilateral covariates, and $\alpha$, $\beta$, $\gamma$ are the coefficient vectors. Commonly used covariates include geographic proximity, the existence of trade agreements, colonial ties, tariffs, non-tariff barriers, or shared languages [11]. The structural gravity model corrects eq. (1) with import and export multilateral resistance terms, which account for the relative nature of bilateral trade shares. This adjustment has been shown to align with various microeconomic models [7]. Gravity models have been widely used to study agrifood trade. For instance, [12] estimate residual trade costs based on a micro-founded gravity equation, finding ad valorem costs to be 60% higher in the global South compared to the North. Studies have also investigated the impact of global and regional trade agreements [13, 14] and the effect of eliminating tariffs [15, 16].
@@
The gravity-based approach is attractive to researchers due to its interpretability, mathematical simplicity, and consistency with various microeconomic theories [9]. However, it is not without its limitations. For one, multilateral trade resistance terms, central to the structural gravity model, are unobservable and must be estimated, often using fixed effects. Elasticity and other key parameters are often unavailable at a granular level, requiring aggregation that can introduce bias [17]. The model’s cost function also depends heavily on the choice of covariates and functional form, making specification crucial for interpreting results. In addition, unobservables—such as the subtle effects of changing political relations, public preferences, or aversions toward products from specific countries—are absorbed in the error term. Finally, while trade costs are generally asymmetric ($C_{ij}\neq C_{ji}$), commonly used covariates are not, making it difficult for a model to capture the inherent imbalances in trade relationships. See [11, 9, 18] for a deeper discussion of challenges and best practices.
@@
In this work, we present a more general approach that dispenses with the use of covariates and a functional form, instead inferring the cost directly from data. Our method is based on the optimal transport (OT) framework [19], which generalises gravity-based models. In OT, trade flows are assumed to match supply and demand to minimise an overall cost. Mathematically, this is expressed as follows: let $\mathbf{C}\in\mathbb{R}^{m\times n}$ be a matrix quantifying the ‘cost’ (in a general sense) of moving goods from country $i$ to $j$. Given the supply vector $\bm{\mu}\in\mathbb{R}^{m}$ and the demand vector $\bm{\nu}\in\mathbb{R}^{n}$, the optimal transport problem consists in finding a transport plan, i.e. a matrix $\mathbf{T}\in\mathbb{R}^{m\times n}_{+}$ with entries $T_{ij}$ modelling the total volume (or value) of transport from country $i$ to $j$, such that the total cost
$c(\mathbf{T})=\sum_{i,j}T_{ij}C_{ij}$ (3)
@@
$c_{\epsilon}(\mathbf{T})=c(\mathbf{T})+\epsilon\mathcal{H}(\mathbf{T}),$ (5)
@@
where $\mathcal{H}(\mathbf{T})$ denotes the entropy of $\mathbf{T}$ and $\epsilon>0$ is a regularisation parameter. It can be shown that the solution will then be of the form
$\mathbf{T}=\bm{\Pi}e^{-\mathbf{C}/\epsilon}\bm{\Omega},$ (6)
where $\bm{\Pi}$ and $\bm{\Omega}$ are diagonal scaling matrices which ensure that the marginal constraints hold (see Methods).
@@
As described in [20], gravity models can be reformulated as solutions of a regularised OT problem with an appropriate choice of parameters. While OT-based models might appear to suggest a centralised control of flows, its dual formulation admits an alternative, decentralised interpretation (see Methods). Indeed, the dual problem can be interpreted as importers seeking to minimise the cost of purchasing commodities and exporters seeking to maximise their profit. The solution at equilibrium coincides with the solution of the OT problem [21], which in its classic form (3)–(4) is well understood. This is less true for the corresponding inverse problem we are interested in, despite its mathematical and practical importance: given a (possibly noisy) observation of $\mathbf{T}$, $\bm{\mu}$ and $\bm{\nu}$, this problem consists in inferring the underlying cost $\mathbf{C}$. As shown in [22], maximum likelihood estimation of underlying gravity model parameters can again be reformulated as an inverse optimal transport problem.
@@
A Ukrainian wheat exports in metric tons, 2021 (left) and 2022 [Uncaptioned image] [Uncaptioned image] C Change in trade volume and cost, selected countries [Uncaptioned image] : Figure 1Ukrainian wheat exports, 2021–2022. A Network of Ukrainian exports, 2021 and 2022. Shown are the largest trading partners, making up 99% of Ukrainian exports. The blue node represents the total Ukrainian export volume (in metric tons), the red nodes are the import volumes. Edge widths represent the flow volume. B The change in trade volume (left) and trade cost (right) for the largest trading partners. C Percent change in trade volume (left bar) and change in trade cost (right bar) for selected countries.
@@
The inference methodology presented in this work is a novel deep learning approach to solve the inverse OT problem, based on recent work on neural parameter calibration [23, 24]. We assume no underlying covariate structure, but instead infer a general cost matrix $\mathbf{C}$, parametrized as a deep neural network, directly from data. We train a neural network $u$ to recognise cost matrices from observations of transport plans for the global food and agricultural trade from 2000–2022 (the ‘training data’) by constraining it to satisfy eq. (6). Put simply, this means fitting the mathematical optimal transport equation to the data in such a way that the predicted cost matrices $\mathbf{C}(t)$ reproduce the observations $\mathbf{T}(t)$. The trained neural network then solves the inverse problem
$\mathbf{C}(t)=u(\mathbf{T}(t))$ (7)
@@
on the observations. Though its ability to extrapolate to new observations depends on the amount of training data, its performance on the training data itself does not. A probability density $\rho$ on the estimates is then naturally obtained by ‘pushing’ the uncertainty on $\mathbf{T}$ through $u$, i.e.
$\rho(\mathbf{C})=u(\rho(\mathbf{T}))$ (8)
@@
(see Methods). As we demonstrate, this approach produces trade flow estimates that are orders of magnitude more accurate than those of a covariate-based gravity model.
@@
The dataset under consideration was assembled by the Food and Agricultural Organisation of the United Nations (FAO), which provides global trade matrices for over 500 products on its portal [^1] https://www.fao.org/faostat/en/#home [25]. Though extensive, many entries in the trade matrices are missing. Furthermore, the FAO reports two values for each bilateral flow $T_{ij}$: one reported by the exporter, and one reported by the importer. There is often a considerable discrepancy between the two, due to a multitude of epistemic factors the FAO lists in its accompanying report [^2] https://files-faostat.fao.org/production/TM/TM_e.pdf. The uncertainty on our estimates naturally follows the uncertainty on the FAO data, without presupposing an underlying statistical model.
@@
We apply our method to analyse global commodity flows from 2000–2022, examining the impacts of events, conflicts, trade agreements, and political changes on trade. The cost matrix uncovers economic effects that are not evident in trade volumes or retail prices alone. The article begins with a study of the war in Ukraine’s impact on global wheat trade, followed by an analysis of free trade agreements and disputes in the Asia-Pacific, as well as the United Kingdom’s 2016 exit from the European Union (Brexit). Finally, we compare our approach to a traditional gravity model, demonstrating its superior performance in both prediction accuracy and uncertainty.
@@ Results
A Russian wheat exports in metric tons, 2021 (left) and 2022 [Uncaptioned image] [Uncaptioned image] C Change in trade volume and cost, selected countries [Uncaptioned image] : Figure 2The same plots as in figure LABEL:fig:Ukraine_wheat with Russia as the exporting partner.
@@
The Russian Federation’s invasion of Ukraine in early 2022 sent shock waves through global food markets [26]. Russia and Ukraine are two of the largest exporters of wheat, together accounting for almost 28% of global wheat exports in 2020. The blockade of trading routes through the Black Sea and the closure or destruction of ports in Mykolaiv and Kherson meant a drop in trade to the overwhelming majority of Ukraine’s export destinations, in some cases by as much as 100% (fig. LABEL:fig:Ukraine_wheatA–B). An increase of wheat exports only occurred to Europe, most significantly to Poland, Spain, Slovakia, and Romania, as well as slight increases to Algeria, India, and Türkiye. However, our analysis shows that, although trade shrank across the globe, the accompanying increase in trade costs disproportionately affected the global South, in particular African nations. Of the ten countries with the largest rise in import costs, four are in Africa, and all are in the global South, while of the ten countries with the largest decrease in trade barriers with Ukraine, seven are in Europe. Countries such as Tanzania or Tunisia, while experiencing a similar drop in trade as the US or France, simultaneously saw an increase in their trade costs. Canadian imports fell by 75%, yet trade utility remained constant, while similar drops in Syria or Egypt led to marked increases in trading barriers. European countries saw an average -0.22 point drop in trade barriers with Ukraine, while the African continent saw an average 0.03 point increase. Imports of wheat from Russia also fell globally (fig. LABEL:fig:Russia_wheat), again affecting Africa particularly severely. European imports of Russian wheat fell by around 40% with a 0.05 point increase in trade costs; African imports fell by on average 71% with a 0.27 point increase in trade costs. While many European countries saw their imports of Ukrainian wheat rise, Russian imports fell sharply. Two notable exceptions in our model are the United Kingdom and Netherlands, which saw increases in trading costs with Ukraine of 0.22 and 0.32 respectively. The two largest hubs for Russian wheat, Egypt and Türkiye, saw no change in their import volumes or import barriers. Meanwhile, Iran saw a 0.6 point increase in trade barriers, leading to a 98% percent decline in Ukrainian wheat imports. For Russian wheat, the estimated increase in trade barriers was only 0.05, leading to a drop in imports of 46%. Russian-Iranian trade barriers were thus not markedly affected by the war, despite a drop in trade volumes.
@@
A Export of sugar products to China, selected countries [Uncaptioned image] B Australian exports to China, selected commodities [Uncaptioned image] : Figure 3Trade with China. A Export of sugar products to China. Top row: estimated trade volume (light blue) in metric tons, as well as the reported values. Bottom row: estimated cost, together with the ASEAN and non-ASEAN averages. B Australian exports to China, selected commodities. Top row: model estimated flow and FAO data; bottom row: estimated cost. Indicated are the signing of ChAFTA (2015, green dotted line) as well as the start of the US-China and Australia-China trade disputes (2018 and 2020, red dotted lines). Errorbands indicate one standard deviation.
@@
A Global barley trade, 2018 (left) and 2021 [Uncaptioned image] B Barley export cost to China, 2018 (left) and 2021 [Uncaptioned image] [Uncaptioned image] : Figure 4Global barley trade between 2015–2022. After the introduction of Chinese import tariffs on Australian barley in May 2020, the entire supply chain restructured itself, with Chinese demand being supplied from France, Canada, and Ukraine, and Australia increasingly exporting to Saudi Arabia and Southeast Asia. A–B Trade in in metric tons, 2018 and 2021. Import values are shown in red, export values in blue. C Model estimated trade volumes (top row) cost (bottom row) for selected countries. Dotted lines indicate the start of the US-China and Australia-China trade wars.
   A Global soya bean trade, 2016 (left) and 2018 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] : Figure 5Global soya bean trade. A In 2018, the Chinese government raised import tariffs on American soya beans in a retaliatory action against US trade restrictions. The shortfall was met by imports from Brazil. B Soya bean yield in 100 g/hectare. Argentina in 2018 experience a major drop in yields, leading to an increase in exports from the US C Predicted trade volumes in metric tons (top row) and predicted cost (bottom row).
[Uncaptioned image] [Uncaptioned image] : Figure 6Change in UK and Ireland (ROI) imports, 2016–2022. For each exporting country, the left two bars indicate the percent change in trade volume between 2016 and 2022 for the UK and the ROI respectively, the right two bars the change in import costs. A Vegetable imports. Top row: lettuce (including chicory), and other fresh vegetables; middle row: tomatoes; bottom row: cucumbers and gherkins. B Wine imports.
@@
A series of free-trade agreements came into effect in Southeast Asia and the Asia-Pacific region in the 2000s and 2010s, significantly among them the China-Australia Free Trade Agreement (ChAFTA) in 2015, the ASEAN-China free trade agreement (ACFTA, gradually entering into force from 2003) and the Comprehensive and Progressive Agreement for Trans-Pacific Partnership (CPTPP) between 11 counties bordering the Pacific Ocean (2018) [27, 28, 29]. Together with China’s accession to the WTO in 2001 and its rapid economic growth, these trade agreements coincide with some of the largest increases in trade flows in recent history. In figure LABEL:fig:China_tradeA, we show trade flow of sugar and sugar products from Thailand, Malaysia, and India to China, as well the estimated costs. In our model, the cost of importing sugar from Thailand fell consistently from 2000–2022, following a general trend for ASEAN countries (bottom row, green line) which commenced around 2005. Indian exports, by comparison, remained relatively low up until 2015, when Indian prime minister Narendra Modi visited China, and top officials from both sides agreed to increase bilateral trade to US\$100 billion by the end of the year. This visit marked a dramatic shift in Indo-Chinese trade, as exemplified by the huge increase of sugar trade. From 2015–2018, export cost from India dropped sharply by 33%, precipitating a steep increase in trade starting in 2018. By contrast, trade cost from non-ASEAN members has remained constant over the past twenty years (red line, fig. LABEL:fig:China_tradeA).
@@
The PRC is one of Australia’s largest export markets for food and agricultural products. Our analysis shows a precipitous reduction in trade barriers for Australian exports since China’s accession to the WTO in 2001 (see fig. LABEL:fig:China_tradeB), particularly for beef, wheat, wine, and sugar. Between 2002 and 2010, these commodities saw a 30–80% drop in trade barriers. Our estimates indicate that ChAFTA had little effect on Australian trade costs, since it succeeded a period of deepening ties. Dairy barriers, for instance, had already fallen from 0.35 to 0.14 from 2000 to 2015, thereafter falling a further 0.03 points until 2020. Wine exports too saw their largest reductions in trade barriers between 2000 and 2010, only experiencing a 0.04 drop from 2015 to 2018 compared to the 0.59 point reduction from 2000–2015.

### v3

@@ Abstract
Global trade is shaped by a complex mix of factors beyond supply and demand, including tangible variables like transport costs and tariffs, as well as less quantifiable influences such as political and economic relations. Traditionally, economists model trade using gravity models, which rely on explicit covariates that might struggle to capture these subtler drivers of trade. In this work, we employ optimal transport and a deep neural network to learn a time-dependent cost function from data, without imposing a specific functional form. This approach consistently outperforms traditional gravity models in accuracy and has similar performance to three-way gravity models, while providing natural uncertainty quantification. Applying our framework to global food and agricultural trade, we show that the Global South suffered disproportionately from the war in Ukraine’s impact on wheat markets. We also analyse the effects of free-trade agreements and trade disputes with China, as well as Brexit’s impact on British trade with Europe, uncovering hidden patterns that trade volumes alone cannot reveal.
@@
1 Department of Applied Mathematics and Theoretical Physics, University of Cambridge, Cambridge CB3 0WA, United Kingdom; 2 Department of Mathematics, Imperial College London, London SW7 2AZ, United Kingdom; 3 School of Business and Management, Queen Mary University of London, London E1 4NS, United Kingdom; 4 Mathematics Institute, University of Warwick, Coventry CV4 7AL, United Kingdom
@@ Introduction
International trade serves as the backbone of the world economy, distributing goods and connecting markets through global logistics networks. Its dynamics are driven by numerous factors beyond mere supply and demand, such as tariffs, non-tariff policy barriers, political and economic tensions, and disruptions caused by accidents, conflicts, and civil wars. Among all traded commodities, agricultural and food products hold particular interest for policymakers and the general public due to their significant volume, high trade value, and critical role in food security and resilience [Friel_et_al2020, Wood_et_al2023]. Consumer food prices are a product of all the complexly interwoven factors governing trade. However, they do not always directly reflect the ease of doing business between any two countries. For instance, in May 2020, China imposed an 80% tariff on Australian barley, leading to a major restructuring of global supply chains (see fig. LABEL:fig:Barley_trade): Chinese demand was suddenly met from France, Canada, and Argentina, while Australia started exporting surplus barley e.g. to Saudi Arabia. Despite these shifts, for the next five months the global barley price barely budged [GIWA_2020, FRED_Barley].
Modelling global trade has garnered significant attention in the economics literature, with gravity models being the most widely used approach [Eaton_Kortum_2002, Anderson_Wincoop_2003, Arkolakis_2012, Allen_2014, Yotov2022]. These models, named for their direct analogy to Newton’s law of gravity, assume that the total trade $T_{ij}$ of a given commodity between two countries $i$ and $j$ is proportional to the total output $O_{i}$ of the source country and the total expenditure $E_{j}$ of the destination country, as well as being inversely related to a ‘distance’ between them:
@@
where $\pi_{i}$ and $\chi_{j}$ are exogenous exporter and importer-side regressors [Dekle_2008], which can be specified as fixed effects, $\rho_{ij}$ are bilateral covariates, and $\alpha$, $\beta$, $\gamma$ are the coefficient vectors. Commonly used covariates include geographic proximity, the existence of trade agreements, colonial ties, tariffs, non-tariff barriers, or shared languages [Yotov_2017]. The structural gravity model corrects eq. (1) with import and export multilateral resistance terms, which account for the relative nature of bilateral trade shares. This adjustment has been shown to align with various microeconomic models [Arkolakis_2012]. Gravity models have been widely used to study agrifood trade. For instance, [Olper_Raimondi2009] estimate residual trade costs based on a micro-founded gravity equation, finding ad-valorem costs to be 60% higher in the Global South compared to the North. Studies have also investigated the impact of global and regional trade agreements [Sarker_Jayasinghe2007, Mujahid_Kalkuhl2016] and the effect of eliminating tariffs [Raimondi_Olper2011, Philippidis_etal2013].
@@
The gravity-based approach is attractive to researchers due to its interpretability, mathematical simplicity, and consistency with various microeconomic theories [Yotov2022]. However, it is not without its limitations. For one, multilateral trade resistance terms, central to the structural gravity model, are unobservable and must be estimated, often using fixed effects [Yotov_2017]. Elasticity and other key parameters are often unavailable at a granular level, requiring aggregation that can introduce bias [Breinlich2022]. The model’s cost function also depends heavily on the choice of covariates and functional form, making specification crucial for interpreting results. In addition, unobservables—such as the subtle effects of changing political relations, public preferences, or aversions toward products from specific countries—are absorbed in the error term. Finally, while trade costs are generally asymmetric ($C_{ij}\neq C_{ji}$), commonly used covariates are not, making it difficult for a model to capture the inherent imbalances in trade relationships. See [Yotov_2017, Yotov2022, Capoani2023] for a deeper discussion of challenges and best practices.
@@
In this work, we present a more general approach that dispenses with the use of covariates and a functional form, instead inferring the cost directly from data. Our method is based on the optimal transport (OT) framework [Villani_2021], which generalises gravity-based models. In OT, trade flows are assumed to match supply and demand to minimise an overall cost. Mathematically, this is expressed as follows: let $\bm{C}\in\mathbb{R}^{m\times n}$ be a matrix quantifying the ‘cost’ (in a general sense) of moving goods from country $i$ to $j$. Given the supply vector $\bm{\mu}\in\mathbb{R}^{m}$ and the demand vector $\bm{\nu}\in\mathbb{R}^{n}$, the optimal transport problem consists in finding a transport plan, i.e. a matrix $\bm{T}\in\mathbb{R}^{m\times n}_{+}$ with entries $T_{ij}$ modelling the total volume (or value) of transport from country $i$ to $j$, such that the total cost
$c(\bm{T})=\sum_{i,j}T_{ij}C_{ij}$ (3)
@@
$c_{\varepsilon}(\bm{T})=c(\bm{T}){-}\varepsilon\mathcal{H}(\bm{T}),$ (5)
@@
where $\mathcal{H}(\bm{T})=-\sum_{ij}T_{ij}(\log T_{ij}-1)$ denotes the negative entropy of $\bm{T}$ and $\varepsilon>0$ is a regularisation parameter. It can be shown that the solution will then be of the form
$\bm{T}=\bm{\Pi}e^{-\bm{C}/\varepsilon}\bm{\Omega},$ (6)
@@
where $\bm{\Pi}$ and $\bm{\Omega}$ are diagonal scaling matrices which ensure that the marginal constraints hold (see Methods). As described in [Wilson_1967], gravity models can be reformulated as solutions of a regularised OT problem with an appropriate choice of parameters. While OT-based models might appear to suggest a centralised control of flows, its dual formulation admits an alternative, decentralised interpretation of importers seeking to minimise the cost of purchasing commodities and exporters seeking to maximise their profit (see Methods). The solution at equilibrium coincides with the solution of the OT problem [Galichon_2016], which in its classic form (3)–(4) is well understood. This is less true for the corresponding inverse problem we are interested in, despite its mathematical and practical importance: given a (possibly noisy) observation of $\bm{T}$, $\bm{\mu}$, and $\bm{\nu}$, this problem consists in inferring the underlying cost $\bm{C}$. Maximum likelihood estimation of costs relates to the inverse optimal transport problem; however, the specific parametrization of the cost matrix in gravity models demands careful estimator design. Moreover, zeros and heteroscedasticity in observed trade flows cause misspecification in gravity model estimation, affecting estimator performance (see [silva2006log]).
@@
A Ukrainian wheat exports in metric tons, 2021 (left) and 2022 [Uncaptioned image] B Percent change in trade volume (left) and absolute change in cost [Uncaptioned image] C Change in trade volume and cost, selected countries [Uncaptioned image] : Figure 1Ukrainian wheat exports, 2021–2022. A Network of Ukrainian exports, 2021 and 2022. Shown are the largest trading partners, making up 99% of Ukrainian exports. The blue node represents the total Ukrainian export volume (in metric tons), the red nodes are the import volumes. Edge widths represent the flow volume. B The change in trade volume (left) and trade cost (right) for the largest trading partners. The large relative increase in trade to India is small in absolute terms and statistically not significant. C Percent change in trade volume (left bar) and change in trade cost (right bar) for selected countries.
@@
The inference methodology presented in this work is a novel deep learning approach to solve the inverse OT problem, based on recent work on neural parameter calibration [Gaskin_2023, Gaskin_2024]. We assume no underlying covariate structure of trade costs, but instead infer a general cost matrix $\bm{C}$, parametrized as a deep neural network, directly from data on trade flows. We train a neural network $u$ to recognise cost matrices from observations of transport plans for the global food and agricultural trade from 2000–2022 (the ‘training data’) by constraining it to satisfy eq. (6). Put simply, this means fitting the mathematical optimal transport equation to the data in such a way that the predicted cost matrices $\bm{C}(t)$ reproduce the observations $\bm{T}(t)$. The trained neural network then solves the inverse problem
$\bm{C}(t)=u(\bm{T}(t))$ (7)
@@
on the observations. Though its ability to generalise to new observations depends on the amount of training data, its performance on the training data itself does not. A probability density $\rho_{C}$ on the estimates is then naturally obtained as the pushforward measure
$\rho_{C}=u_{\#}\rho_{T},$ (8)
@@
where $\rho_{T}$ is the measure on $\bm{T}$. Additionally, we train a family of neural networks to capture the spread in cost matrices that optimally reproduce the transport plan (see Methods). As we demonstrate, this approach produces trade flow estimates that are an order of magnitude more accurate than those of a traditional covariate-based gravity model.
@@
The dataset under consideration was assembled by the Food and Agricultural Organisation of the United Nations (FAO), which provides global trade matrices for over 500 products on its portal111fao.org/faostat/en/#home [FAOStat]. Though extensive, many entries in the trade matrices are missing. Furthermore, the FAO reports two values for each bilateral flow $T_{ij}$: one reported by the exporter, and one reported by the importer. There is often a considerable discrepancy between the two, due to a multitude of epistemic factors the FAO lists in its accompanying report222files-faostat.fao.org/production/TM/TM_e.pdf. The uncertainty on our estimates naturally follows the uncertainty on the FAO data, without presupposing an underlying statistical model.
@@
We apply our method to analyse global commodity flows from 2000–2022, examining the impacts of events, conflicts, trade agreements, and political changes on trade. The cost matrix uncovers economic effects that are not evident in trade volumes or retail prices alone. The article begins with a study of the war in Ukraine’s impact on global wheat trade, followed by an analysis of free trade agreements and disputes in the Asia-Pacific, as well as the United Kingdom’s 2016 exit from the European Union (Brexit). We demonstrate our method’s ability to provide meaningful uncertainty quantification, and compare it to a traditional gravity model, demonstrating superior prediction accuracy.
@@ Results
The Russian Federation’s invasion of Ukraine in early 2022 sent shock waves through global food markets [Laber_2023]. Russia and Ukraine are two of the largest exporters of wheat, together accounting for almost 28% of global wheat exports in 2020. The blockade of trading routes through the Black Sea and the closure or destruction of ports in Mykolaiv and Kherson meant a drop in trade to the overwhelming majority of Ukraine’s export destinations, in some cases by as much as 100% (fig. LABEL:fig:Ukraine_wheatAB). An increase of wheat exports only occurred to Europe, most significantly to Poland, Spain, Slovakia, Romania, as well as to Algeria, and Türkiye. However, our analysis shows that, although trade shrank across the globe, the accompanying increase in wheat trade costs disproportionately affected the Global South, in particular African nations. Of the ten countries with the largest rise in wheat import costs, five are in Africa, and all are in the Global South, while of the ten countries with the largest decrease in trade barriers with Ukraine, nine are in Europe (see figs. LABEL:fig:SI:Wheat_macrotrends_1, LABEL:fig:SI:Wheat_macrotrends_2, and LABEL:fig:SI:Wheat_macrotrends_3 in the SI). Countries such as Nigeria or Angola, while experiencing a similar drop in trade as the Norway or France, simultaneously saw an increase in their trade costs. Canadian imports fell by 91%, yet unit trade costs remained nearly constant, while similar drops in Zambia or the DR Congo led to marked increases in unit trade costs, pointing to trading barriers. European countries saw an average 9% drop in unit wheat trade costs with Ukraine, while Sub-Saharan Africa saw an average 22% increase (see fig. LABEL:fig:SI:Wheat_macrotrends_2 in the SI). Imports of wheat from Russia also fell globally (see fig. LABEL:fig:SI:Russia_wheat in the SI), again affecting Africa particularly severely. European imports of Russian wheat fell by around 74% with an 18% increase in trade costs; African imports fell by on average 80% with a 36% increase in trade costs. While many European countries saw their imports of Ukrainian wheat rise, Russian imports fell sharply. The two largest hubs for Russian wheat, Egypt and Türkiye, saw no change in their import volumes and small declines in their import barriers. Meanwhile, Iran saw a 0.51 point increase in wheat unit trade costs, leading to a 97% percent decline in Ukrainian wheat imports. For Russian wheat, the estimated increase in trade costs was only 0.04, leading to a drop in imports of 53%. Russian-Iranian trade barriers were thus not markedly affected by the war, despite a drop in trade volumes.
@@
A Export of sugar products to China, selected countries [Uncaptioned image] B Australian exports to China, selected commodities [Uncaptioned image] : Figure 2Trade with China. A Export of sugar products to China. Top row: estimated trade volume (light blue) in metric tons, as well as the reported values. Bottom row: estimated cost, together with the ASEAN and non-ASEAN averages. B Australian exports to China, selected commodities. Top row: model estimated flow and FAO data; bottom row: estimated cost. Indicated are the signing of ChAFTA (2015, green dotted line) as well as the start of the US-China and Australia-China trade disputes (2018 and $20$, red dotted lines). Errorbands indicate one standard deviation.
@@
A Global barley trade, 2018 (left) and 2021 [Uncaptioned image] B Barley export cost to China, 2018 (left) and 2021 [Uncaptioned image] C Trade volumes and costs [Uncaptioned image] : Figure 3Global barley trade between 2015–2022. After the introduction of Chinese import tariffs on Australian barley in May 2020, the entire supply chain restructured itself, with Chinese demand being supplied from France, Canada, and Ukraine, and Australia increasingly exporting to Saudi Arabia and Southeast Asia. A–B Trade in in metric tons, 2018 and 2021. Import values are shown in red, export values in blue. C Model estimated trade volumes (top row) and cost (bottom row) for selected countries. Dotted lines indicate the start of the US-China and Australia-China trade wars.
A Global soya bean trade, 2016 (left) and 2018 [Uncaptioned image] B Yield [Uncaptioned image] C Trade volumes and utilities [Uncaptioned image] : Figure 4Global soya bean trade. A In 2018, the Chinese government raised import tariffs on American soya beans in a retaliatory action against US trade restrictions. The shortfall was met by imports from Brazil. B Soya bean yield in 100 g/hectare. Argentina in 2018 experienced a major drop in yields, leading to an increase in exports from the US. C Predicted trade volumes in metric tons (top row) and predicted cost (bottom row).
A Vegetables [Uncaptioned image] B Wine [Uncaptioned image] : Figure 5Change in UK and Ireland (ROI) imports, 2016–2022. For each exporting country, the left two bars indicate the percent change in trade volume between 2016 and 2022 for the UK and the ROI respectively, the right two bars show the change in import costs. A Vegetable imports. Top row: lettuce (including chicory), and other fresh vegetables; middle row: tomatoes; bottom row: cucumbers and gherkins. B Wine imports.
@@
A series of free-trade agreements came into effect in Southeast Asia and the Asia-Pacific region in the 2000s and 2010s, significantly among them the China-Australia Free Trade Agreement (ChAFTA) in 2015, the ASEAN-China free trade agreement (ACFTA, gradually entering into force from 2003) and the Comprehensive and Progressive Agreement for Trans-Pacific Partnership (CPTPP) between 11 counties bordering the Pacific Ocean (2018) [ChAFTA, ACFTA, CPTPP]. Together with China’s accession to the WTO in 2001 and its rapid economic growth, these trade agreements coincide with some of the largest increases in trade flows in recent history. In figure LABEL:fig:China_tradeA, we show the trade flow of sugar and sugar products from Thailand, Malaysia, and India to China, as well as the estimated costs. In our model, the cost of importing sugar from Thailand fell consistently from 2000–2022, following a general trend for ASEAN countries (bottom row, green line) which commenced around 2005. Indian exports, by comparison, remained relatively low until 2015, when Indian prime minister Narendra Modi visited China, and top officials from both sides agreed to increase bilateral trade to US\$100 billion by the end of the year. This visit marked a dramatic shift in Indo-Chinese trade, as exemplified by the huge increase in sugar trade. From 2015–2022, sugar export cost from India dropped sharply by 33%, precipitating a steep increase in trade. By contrast, trade cost from non-ASEAN members has remained constant over the past twenty years (red line, fig. LABEL:fig:China_tradeA).
@@
The PRC is one of Australia’s largest export markets for food and agricultural products. Our analysis shows a precipitous reduction in trade barriers for Australian exports since China’s accession to the WTO in 2001 (see fig. LABEL:fig:China_tradeB), particularly for beef, wheat, wine, and dairy. Between 2002 and 2010, these commodities saw a 30–50% drop in their respective trade costs. Our estimates indicate that ChAFTA had little effect on Australian trade costs, since it succeeded a period of deepening ties. Dairy trade costs, for instance, had already fallen from 0.51 to 0.1 from 2000 to 2015, thereafter only falling a further 0.04 points until 2020. Wine exports too saw their largest reductions in trade costs between 2000 and 2010, only experiencing a 0.07 drop from 2015 to 2018 compared to the 0.52 point reduction from 2000–2015.